OpenStudy (anonymous):

What is the solution to the equation? 7(5 + 9) = 7 · x + 7 · 9 A. x = 1 B. x = 5 C. x = 7 D. x = 9 @iGreen @iGreen. @sleepyjess

OpenStudy (sleepyjess):

Substitute in the options for x and see which one works out.

OpenStudy (dtan5457):

Well the first part of the equation 7(5+9)=98. For the 2nd part, 7 times 9 is 63, 98-63=35 7 times what is 35?

OpenStudy (sleepyjess):

OR 7(5+9)=7*5+7*9 35+63=7*x+7*9 98=7x+63

OpenStudy (dtan5457):

In case you want to check your answer it would be B. x = 5

OpenStudy (anonymous):

7*9 from both sides cancels out and simply 7x=7*5 or x=5